# How to Build a Quantum Supercomputer: Scaling Challenges and Opportunities

### v1

@@ B.3 Comparison with the AzureQRE toolkit
Baseline Parameter Set ($\bm{\Lambda\approx 2.34}$) Target Parameter Set ($\bm{\Lambda\approx 9.3}$) Desired Parameter Set ($\bm{\Lambda\approx 18}$)
@@
$1.0$ mHa 6 - - - 1.2$\times$107 28.8 minutes 77 4.1$\times$106 16.1 minutes 43
18 - - - 2.4$\times$107 1.2 days 91 7.8$\times$106 15.7 hours 51
26 - - - 3.3$\times$107 4.8 days 95 1.0$\times$107 2.7 days 53
76 - - - 1.0$\times$108 1.4 years 111 3.2$\times$107 289.0 days 63
@@
$1.0$ mHa 6 1.5$\times$107 5.4 minutes 25, 61 | 75 1.6$\times$106 1.2 minutes 23 | 27 8.9$\times$105 57.6 seconds 17 | 21
18 4.9$\times$107 2.2 days 15, 33, 75 | 91 4.0$\times$106 12.4 hours 11, 29 | 33 2.1$\times$106 10.2 hours 21 | 27
26 6.0$\times$107 10.0 days 15, 33, 87 | 103 5.5$\times$106 2.3 days 11, 31 | 37 2.8$\times$106 1.7 days 23 | 27
76 1.2$\times$108 2.6 years 17, 37, 93 | 109 1.5$\times$107 234.1 days 13, 33 | 43 8.0$\times$106 168.8 days 27 | 31
@@ Appendix C Double-factorized quantum chemistry
The fourth–order Coulomb tensor $h_{ijkl}$ can be written as a $N_{o}^{2}/4\times N_{o}^{2}/4$ electronic repulsion integral (ERI) matrix, $A$, where $N_{o}$ is the number of spin orbitals. $A$ is positive semi-definite and generally has a rank $L=\mathcal{O}(N)$ for chemical systems. We can diagonalize $A$, leading to a decomposition in terms of an auxiliary tensor $\mathcal{L}$ such that Motta et al. (2021b):
@@
where $\tilde{h}_{ij}\equiv h_{ij}-\frac{1}{2}\sum_{l}h_{illj}$ comes from the reordering of the creation and annihilation operators. By truncating some of the eigenvalues, a low–rank approximation can be obtained. There exists efficient walk operators, $W$, which implement this Hamiltonian as a quantum circuit, as described in Ref. von Burg et al. (2021b).
@@ Appendix D Runtime of classical algorithms for quantum chemistry
To provide realistic estimates of the classical resources required for various classical quantum chemistry algorithms, we extrapolate the results of recent publications that use full configuration interaction (FCI) Gao et al. (2024) and density matrix renormalization group (DMRG) Bellonzi et al. (2024). For the FCI calculations, we note that Ref. Gao et al. (2024) reports running their largest system of C3H8 in an STO-3G basis, which has 26 electrons in 23 orbitals, a calculation involving 1.3 trillion determinants, took 113.6 hours using 512 processes, a total of around 58k CPU hours. Assuming quadratic scaling with number of determinants ($\mathcal{O}(N_{det}^{2})$) scaling for the FCI algorithm, we use this single data point to compute a realistic prefactor for the computational time scaling. Note that the number of determinants scales exponentially with number of orbitals. We then take the worst-case number of determinants for each number of orbitals, where the number of electrons ($N_{e}$) is equal to the number of orbitals ($N_{o}$), and calculate the total number of determinants as $N_{det}=\big{(}N_{o}!/(N_{o}-N_{e})!N_{e}!\big{)}^{2}$ and, assuming a factor of 1000 in parallelism, compute the time for various numbers of orbitals, consistent with the 512 CPUs used in Ref. Gao et al. (2024).
For the DMRG calculations, we assume cubic scaling with bond dimension ($\mathcal{O}(\chi^{3})$). Note that there is no definitive scaling of bond dimension $\chi$ with number of orbitals for generic quantum chemistry problems, but it is generally expected to scale exponentially for strongly correlated systems. Ref. Bellonzi et al. (2024) reports estimates of the necessary bond dimension for various homogeneous catalysts as well as runtimes for smaller bond dimension DMRG calculations. Using the data in Table 3 of Ref. Bellonzi et al. (2024), specifically the data which was run on a computer cluster, we fit parameters $a$ and $b$ in the scaling function $f(\chi)=a\chi^{3}+b$ and then use those coefficients to predict the runtime necessary for the reported bond dimensions necessary to reach chemical accuracy, assuming a factor of 100 parallelism, consistent with the 40 CPUs used in the Ref Bellonzi et al. (2024).
@@ Appendix E Circuit-level noise model
We employ a circuit-level depolarizing noise model for the benchmarking and sensitivity analysis simulations in Section III.2. It consists of three types of errors: gate errors, idling errors, and state preparation and measurement (SPAM) errors, where the strength of each type is determined from the hardware noise parameters, such as those found in Table 1.
Imperfect gates are modeled by adding a depolarizing noise channel at rate $p$ to the gate qubits after the application of each gate. For one-qubit gates, the noise channel randomly applies one of $X$, $Y$ or $Z$, each with probability $p/3$. Similarly, for two-qubit gates, the noise channel applies one of the 15 two-qubit non-identity Pauli gates, each with probability $p/15$. The rate $p$ is determined by utilizing the formula for the depolarizing channel’s average gate fidelity,
@@
Any qubit that is idling experiences an error, dependent on both the decoherence time $T_{1}$ of the qubit and the time $t$ it takes to apply the gate(s) to the active qubits. The error is modeled with a single-qubit depolarizing noise channel with rate equal to
@@
The errors in state preparation and reset are captured by assuming that with rate $p$ the orthogonal state is produced, i.e., $|0\rangle$ is prepared instead of $|1\rangle$ and vice versa. Similarly, measurements in the $Z$ basis are flipped at rate $p$. In all three cases the fidelity of the operation is
$F_{\text{SPAM}}=\frac{P(0|0)+P(1|1)}{2},$ (46)
from which we directly determine the rate $p=1-F_{\text{SPAM}}$.

### v2

@@ B.3 Comparison with the AzureQRE toolkit
Baseline Parameter Set Target Parameter Set Desired Parameter Set
@@
$1.6$ mHa 6 - - - 1.1$\times$107 17.4 minutes 75 3.1$\times$106 10.0 minutes 43
18 - - - 2.3$\times$107 16.2 hours 89 7.6$\times$106 9.3 hours 51
26 - - - 3.1$\times$107 2.8 days 93 9.9$\times$106 1.6 days 53
76 - - - 9.7$\times$107 286.3 days 109 3.0$\times$107 160.2 days 61
@@
$1.6$ mHa 6 1.4$\times$107 3.3 minutes 25, 59 | 73 1.3$\times$106 1.2 minutes 23 | 27 8.2$\times$105 35.8 seconds 17 | 21
18 4.8$\times$107 1.4 days 15, 31, 75 | 93 3.6$\times$106 7.3 hours 11, 27 | 33 1.7$\times$106 5.6 hours 21 | 25
26 5.5$\times$107 5.4 days 15, 33, 79 | 95 4.7$\times$106 1.3 days 11, 29 | 35 2.6$\times$106 23.6 hours 23 | 27
76 1.2$\times$108 1.4 years 17, 37, 89 | 107 1.3$\times$107 121.8 days 13, 33 | 39 7.7$\times$106 96.8 days 25 | 31
@@ Appendix C Double-factorized quantum chemistry
The fourth-order Coulomb tensor $h_{ijkl}$ can be written as a $N_{o}^{2}/4\times N_{o}^{2}/4$ electronic repulsion integral (ERI) matrix, $A$, where $N_{o}$ is the number of spin orbitals. $A$ is positive semi-definite and generally has a rank $L=\mathcal{O}(N)$ for chemical systems. We can diagonalize $A$, leading to a decomposition in terms of an auxiliary tensor $\mathcal{L}$ such that Motta et al. (2021b):
@@
where $\tilde{h}_{ij}\equiv h_{ij}-\frac{1}{2}\sum_{l}h_{illj}$ comes from the reordering of the creation and annihilation operators. By truncating some of the eigenvalues, a low–rank approximation can be obtained. There exist efficient walk operators, $W$, which implement this Hamiltonian as a quantum circuit, as described in Ref. von Burg et al. (2021b).
@@ Appendix D Runtime of classical algorithms for quantum chemistry
To provide realistic estimates of the classical resources required for various classical quantum chemistry algorithms, we extrapolate the results of recent publications that use full configuration interaction (FCI) Gao et al. (2024) and density matrix renormalization group (DMRG) Bellonzi et al. (2024a). For the FCI calculations, we note that Ref. Gao et al. (2024) reports running their largest system of C3H8 in an STO-3G basis, which has 26 electrons in 23 orbitals, a calculation involving 1.3 trillion determinants, took 113.6 hours using 512 processes, a total of around 58k CPU hours. Assuming quadratic scaling with number of determinants ($\mathcal{O}(N_{\text{det}}^{2})$) scaling for the FCI algorithm, we use this single data point to compute a realistic prefactor for the computational time scaling. Note that the number of determinants scales exponentially with number of orbitals. We then take the worst-case number of determinants for each number of orbitals, where the number of electrons ($N_{e}$) is equal to the number of orbitals ($N_{o}$), and calculate the total number of determinants as $N_{\text{det}}=\big{(}N_{o}!/(N_{o}-N_{e})!N_{e}!\big{)}^{2}$ and, assuming a factor of 1000 in parallelism, compute the time for various numbers of orbitals, consistent with the 512 CPUs used in Ref. Gao et al. (2024).
For the DMRG calculations, we assume cubic scaling with bond dimension ($\mathcal{O}(\chi^{3})$). Note that there is no definitive scaling of bond dimension $\chi$ with number of orbitals for generic quantum chemistry problems, but it is generally expected to scale exponentially for strongly correlated systems. Estimates of the necessary bond dimension for various homogeneous catalysts are reported in Ref. Bellonzi et al. (2024a), as well as runtimes for smaller bond dimension DMRG calculations. Using the data in Table 3 of Ref. Bellonzi et al. (2024a), specifically the data which was run on a computer cluster, we fit parameters $a$ and $b$ in the scaling function $f(\chi)=a\chi^{3}+b$ and then use those coefficients to predict the runtime necessary for the reported bond dimensions necessary to reach chemical accuracy, assuming a factor of 100 parallelism, consistent with the 40 CPUs used in the Ref. Bellonzi et al. (2024a).
@@ Appendix E Circuit-level noise model
We employ a circuit-level depolarizing noise model for the benchmarking and sensitivity analysis simulations in Section III.2. It consists of three types of errors: gate errors, idling errors, and state preparation and measurement (SPAM) errors, where the strength of each type is determined from the values of hardware noise parameters, such as those provided in Table 1.
Imperfect gates are modelled by adding a depolarizing noise channel at rate $p$ to the gate qubits after the application of each gate. For one-qubit gates, the noise channel randomly applies one of $X$, $Y$, or $Z$, each with probability $p/3$. Similarly, for two-qubit gates, the noise channel applies one of the 15 two-qubit non-identity Pauli gates, each with probability $p/15$. The rate $p$ is determined by utilizing the formula for the depolarizing channel’s average gate fidelity,
@@
Idle qubits are succeptible to errors dependent on both the decoherence time $T_{1}$ of the qubit and the time $t$ it takes to apply the gate(s) to the active qubits. The error is modeled with a single-qubit depolarizing noise channel with rate equal to
@@
The errors in state preparation and reset are captured by assuming that with rate $p$ the orthogonal state is produced, that is, $|0\rangle$ is prepared instead of $|1\rangle$ and vice versa. Similarly, measurements in the $Z$ basis are flipped at rate $p$. In all three cases the fidelity of the operation is
$F_{\textsc{spam}}=\frac{P(0|0)+P(1|1)}{2},$ (49)
from which we directly determine the rate $p=1-F_{\textsc{spam}}$.
Appendix F High-performance real-time control and decoding with DGX Quantum
Nvidia DGX Quantum NVIDIA Corporation (2025a), developed in collaboration with Quantum Machines, is a groundbreaking architecture that tightly integrates advanced classical and quantum computational capabilities. It combines Nvidia Grace Hopper superchips with the Quantum Machines OPX1000 control system, achieving microsecond-scale latency between GPU and QPUs. This connection enables latency-critical work like QEC to be accelerated with GPUs. The system is designed to scale with both quantum and classical computing demands; as quantum computers advance, additional DGX Quantum nodes can be incorporated to interface with more QPUs. Moreover, these nodes can be connected by both classical and quantum interconnects to scale up to large accelerated quantum supercomputing systems, as shown in Figure 3. DGX Quantum is also QPU agnostic, enabling all qubit modalities to be integrated within the architecture.
The Nvidia GH200 superchip is a key part of the DGX Quantum integrated system, delivering exceptional performance for hybrid quantum–classical workloads through its heterogeneous architecture. This architecture eliminates bottlenecks in data movement and ensures optimal performance for latency-sensitive quantum–classical interactions, as required in QEC and hybrid algorithm execution. Advanced features of the Hopper GPU, including asynchronous execution engines and thread block reconfiguration, further enhance computational efficiency, ensuring that all resources are fully utilized in the DGX Quantum integrated system. Future DGX Quantum systems may use Grace Blackwell GB200 superchips (see Figure 46). A comparison of Grace Hopper and Blackwell superchips is shown in Table 8, which demonstrates the GB200 superchip’s significantly improved specifications over those of the GH200. It is an important example of improvement because to calculate the future true costs of quantum computing, one needs to factor in the decreasing cost of error correction due to improvement in classical computing infrastructure.
Figure 46: Picture of the Nvidia Grace-Blackwell GB200 superchip. The GB200 contains two Nvidia Blackwell B200 Tensor Core GPUs with an Nvidia Grace CPU, interconnected via a 900 GB/s NVLink-C2C interface, creating a unified memory architecture. This configuration provides a total of 896 GB of memory, combining 384 GB of HBM3e memory on the GPUs and 512 GB of LPDDR5 memory on the CPU.
GH200 Grace Hopper Superchip GB200 Grace Blackwell Superchip
Configuration 1 Grace CPU : 1 Hopper GPU 1 Grace CPU : 2 Blackwell GPUs
FP4 Tensor Core 20 PFLOPS 40 PFLOPS
FP8/FP6 Tensor Core 4 PFLOPS 20 PFLOPS
INT8 Tensor Core 4 POPS 20 POPS
FP16/BF16 Tensor Core 2 PFLOPS 10 PFLOPS
TF32 Tensor Core 1 PFLOPS 5 PFLOPS
FP32 67 TFLOPS 180 TFLOPS
FP64 34 TFLOPS 90 TFLOPS
FP64 Tensor Core 67 TFLOPS 90 TFLOPS
GPU Memory | Bandwidth 96 GB HBM3 | 4 TB/s Up to 384 GB HBM3e | 16 TB/s
NVLink Bandwidth 0.9 TB/s 3.6 TB/s
CPU Core Count 72 Arm Neoverse V2 cores 72 Arm Neoverse V2 cores
CPU Memory | Bandwidth Up to 480GB LPDDR5X | Up to 512 GB/s Up to 480GB LPDDR5X | Up to 512 GB/s
Table 8: Grace Hopper vs. Grace Blackwell superchip specifications.
DGX Quantum also enables many other control mechanisms, such as error mitigation, dynamic circuit programming, and adaptive circuit knitting. By directly writing quantum control data into the GH200’s local memory, latency overhead is eliminated, ensuring that the system can respond to quantum hardware feedback in real time. This feature is particularly valuable for iterative classical–quantum applications. For less latency-critical scenarios, users can queue quantum programs locally and remotely, enabling resource flexibility while maintaining access to DGX Quantum’s high-performance computational capabilities.
Appendix G CUDA-Q: A software platform for heterogeneous quantum–classical computing
DGX Quantum integrated systems are built to run Nvidia CUDA-Q NVIDIA Corporation (2025b), which is an open source programming framework and compiler toolchain designed as an efficient and unified platform for programming hybrid quantum–classical systems. CUDA-Q supports both Python and C++ programming interfaces. Its kernel-based programming capabilities are specifically tailored for hybrid quantum–classical systems. CUDA-Q programs use the concept of a quantum kernel to differentiate between host and quantum device code, with kernels specifying a target for compilation and execution. The CUDA-Q platform also includes the NVQ++ compiler, which supports split compilation by lowering quantum kernels to a multi-level intermediate representation (MLIR) and a quantum intermediate representation (QIR). This ensures the seamless integration of classical and quantum resources needed for accelerating applications in large-scale quantum computing.
CUDA-Q includes specialized libraries designed to streamline the development of quantum algorithms and support advanced research. The CUDA-Q QEC library provides optimized implementations of key quantum error correction primitives (such as GPU-accelerated decoders) and the CUDA-Q Solvers library includes application development primitives (such as prebuilt optimized kernels for the VQE, ADAPT-VQE, and QAOA methods). CUDA-Q also offers GPU-accelerated quantum dynamics simulations, enabling accurate modeling of open quantum systems—a critical capability for the design, characterization, optimization, and scaling of QPUs.
Another key advantage of CUDA-Q is its ability to scale the performance of quantum simulations. Leveraging Nvidia’s multi-GPU acceleration, CUDA-Q can achieve a massive speedup over traditional CPU-based simulations NVIDIA Corporation (2025b), enabling researchers to efficiently simulate large quantum circuits. CUDA-Q supports multiple simulation technique backends such as state vector and tensor networks, allowing users to select from among various simulation methods for their specific problem.
Additionally, CUDA-Q is interoperable with AI software and the CUDA software ecosystem. This ecosystem is a comprehensive suite of tools, libraries, and frameworks designed to harness the computational power of GPUs for diverse applications, ranging from AI and data science to HPC and quantum simulation. It includes components that provide developers with optimized building blocks, scalable programming models, and integration capabilities for hybrid quantum–classical workflows.